OpenStudy (anonymous):

Choose the correct simplification of the expression 6b / a^-9

ganeshie8 (ganeshie8):

use the exponent property : \(\huge \frac{1}{a^{-m}} = a^m\)

OpenStudy (anonymous):

is the answer 6a^9b

ganeshie8 (ganeshie8):

thats right ! nice work :))

OpenStudy (anonymous):

thank you! can you help me with one more question please

ganeshie8 (ganeshie8):

np post :)

OpenStudy (anonymous):

Choose the correct simplification of the expression (−5a5b4c3)^0

ganeshie8 (ganeshie8):

Easy !! anything to the power 0 is 0

ganeshie8 (ganeshie8):

\(\huge (−5a5b4c3)^0 = 0 \)

OpenStudy (anonymous):

ohhh ok thank you very much :))

ganeshie8 (ganeshie8):

np :) yw :DD